OpenStudy (anonymous):

Given Cos theta =B find all other trig values B^2+sin^2 theta =1 square root 1-B^2= sin is this true?

OpenStudy (sgadi):

yes

OpenStudy (anonymous):

therefor i can draw from that that csc theta = 1/ square root 1-B^2 cec = 1/b and tan = b/square root 1-B^2

OpenStudy (amistre64):

cos(t) = b sin(t) = sqrt(1-b^2) tan(t) = sqrt(1-b^2) / b sec is 1/cos sec = 1/b csc is 1/sin csc = 1/(sqrt(1-b^2)) cot is 1/tan cot = b/sqrt(1-b^2)

OpenStudy (amistre64):

so yes, they are true

OpenStudy (anonymous):

thank you!